OpenStudy (anonymous):

Solve the inequality: 4s + 3(2-3s) ≤ 5(2-s)

OpenStudy (anonymous):

4s+6-9s≤10-5s -5s+6≤10-5s 0<4

OpenStudy (anonymous):

4s+6-9s<=10-5s -5s+5s<=10-6 0<=4 which is true for any s

OpenStudy (anonymous):

so, you have inequality that is true for any variable s (real number )

OpenStudy (anonymous):

thanks! helps alot

OpenStudy (anonymous):

:)